OpenStudy (anonymous):

I am trying to prove that (tanx+cotx)/(secx+cscx) is equal to 1/(cosx+sinx) ... can someone show me the steps?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

Just write all the funny expression out as sin/cos and do the algebra, it will just fall out...

OpenStudy (anonymous):

right, i'm at ((sinx/cosx)+(cosx/sinx))/((1/cosx)+(1/sinx)) and as usual the basic algebra is messing me up as to how i "flip" the fractions.....

OpenStudy (anonymous):

((1/cosx)+(1/sinx)) do the addition first and then turn it upside down n multiply.

OpenStudy (anonymous):

Rewrite as: \[\huge\frac{\frac{\sin(x)}{\cos(x)}+\frac{\cos(x)}{\sin(x)}}{\frac{1}{\cos(x)}+\frac{1}{\sin(x)}}\] No get a common denominator in the top and bottom. \[\huge\frac{\frac{\sin^2(x)+\cos^2(x)}{\sin(x)\cos(x)}}{\frac{\sin(x)+\cos(x)}{\sin(x)\cos(x)}}\] Well: sin^2(x)+cos^2(x)=1 So you have: \[\huge\frac{\frac{1}{\sin(x)\cos(x)}}{\frac{\sin(x)+\cos(x)}{\sin(x)\cos(x)}}=\frac{\sin(x)\cos(x)}{\sin(x)\cos(x)(\sin(x)+\cos(x))}\] \[\huge=\frac{1}{\sin(x)+\cos(x)}\]

OpenStudy (anonymous):

hory smokes. ty ty

OpenStudy (anonymous):

I made it big so you could read the fractions, they were kinda jumbled. :P

OpenStudy (anonymous):

you are mighty in the ways of the trig

OpenStudy (anonymous):

Haha, I try :P I've just seen it alot so the approaches come easy. The approach is the most important part. The algebra can be done by computers :P